OpenStudy (anonymous):

find each function value, if it exists. f(t)=√(t^2+1) A.) f(0)= ? (Simplify your answer. Type an exact answer, using radicals as needed.) B.) The square root is not a real number.

OpenStudy (alfie):

Just plug in the zero inside the function. You'll get square root of one, which is one.

OpenStudy (alfie):

1

OpenStudy (anonymous):

f(-8)?

OpenStudy (alfie):

Same stuff. Plug in the number instead of your variable. In this case "t"

OpenStudy (anonymous):

i dont understand

OpenStudy (alfie):

f(x) = x f(1) = 1 f(x) = x + 2 f(1) = 1 + 2

OpenStudy (alfie):

f(x) = 5x-3 f(-2) = 5(-2) - 3

OpenStudy (anonymous):

f(-8)=?